### v2

# Convergent Primal-Dual Plug-and-Play Image Restoration: A General Algorithm and Applications

Abstract
We propose a general deep plug-and-play (PnP) algorithm with a theoretical convergence guarantee. PnP strategies have demonstrated outstanding performance in various image restoration tasks by exploiting the powerful priors underlying Gaussian denoisers. However, existing PnP methods often lack theoretical convergence guarantees under realistic assumptions due to their ad-hoc nature, resulting in inconsistent behavior. Moreover, even when convergence guarantees are provided, they are typically designed for specific settings or require a considerable computational cost in handling non-quadratic data-fidelity terms and additional constraints, which are key components in many image restoration scenarios. To tackle these challenges, we integrate the PnP paradigm with primal-dual splitting (PDS), an efficient proximal splitting methodology for solving a wide range of convex optimization problems, and develop a general convergent PnP framework. Specifically, we establish theoretical conditions for the convergence of the proposed PnP algorithm under a reasonable assumption. Furthermore, we show that the problem solved by the proposed PnP algorithm is not a standard convex optimization problem but a more general monotone inclusion problem, where we provide a mathematical representation of the solution set. Our approach efficiently handles a broad class of image restoration problems with guaranteed theoretical convergence. Numerical experiments on specific image restoration tasks validate the practicality and effectiveness of our theoretical results.
Index Terms:
Image restoration, plug-and-play (PnP) algorithms, primal-dual splitting (PDS), convergence guarantee.
I Introduction
I-A Image Restoration by Proximal Splitting Algorithms
Image restoration is a longstanding and essential problem with diverse applications, ranging from remote sensing, geoscience, and astronomy to biomedical imaging. In general, it is reduced to the inverse problem of estimating an original image from an observed image, which is degraded by some linear observation process and noise contamination (e.g., additive white Gaussian noise and Poisson noise).
Since this inverse problem is often ill-posed or ill-conditioned, a standard approach formulates the restoration task as a convex optimization problem to characterize a desirable solution. In this scheme, we typically minimize the sum of a data-fidelity term with respect to a given observed image and a regularization term reflecting prior knowledge on natural images. Typically, the data-fidelity is chosen as the negative log-likelihood of the statistical distribution assumed for the noise contamination. Conversely, the choice of the regularization term is highly diverse. One of the most popular techniques is Total Variation (TV) [1] and its extensions[2, 3, 4, 5], which promote the spatial piecewise smoothness of images. In any case, these data-fidelity and regularization terms are possibly nonsmooth, which makes it difficult to solve the convex optimization problem analytically or using gradient-based methods.
This difficulty has been overcome by proximal splitting algorithms, which assume that the proximal operators associated with these terms can be computed efficiently [6, 7, 8]. Notable examples include Forward-Backward Splitting (FBS) [9], Alternating Direction Method of Multipliers (ADMM) [10], and Primal-Dual Splitting (PDS) [11, 12, 13, 14]. Among these algorithms, PDS offers significant flexibility and efficiency, as it can handle nonsmooth convex terms involving linear operators without requiring matrix inversions or performing inner iterations, which is generally impossible in the case of FBS or ADMM. These features offer several advantages for image restoration applications [15, 16] (refer to Section III-B for more detailed explanation with specific formulations).
I-B Plug-and-Play (PnP) Algorithms
TABLE I: Pros and Cons of Existing PnP Methods and the Proposed Method.
Algorithms Authors Data-fidelity
Additional
constraints
Inversion
free
Convergence
guarantees
Solution
characterization
Assumption
on denoiser $J$
PnP-FBS Pesquet et.al. [17] Quadric -- $J$ is firmly nonexpansive.
Ebner et al. [18] Quadric -- $J$ is contractive.
PnP-ADMM Venkatakrishnan et al. [19] $**{}^{\text{*}}$- -- -- -- --
Rond et al. [20] GKL div. -- -- -- -- --
Sreehari et al. [21] $**{}^{\text{*}}$- -- -- $\nabla J$ is a symmetric matrix, etc.
S. H. Chan et al. [22] $**{}^{\text{*}}$- -- -- $J$ is bounded, etc.
P. Nair et al. [23] $**{}^{\text{*}}$- -- -- $J$ is firmly nonexpansive.
Sun et al. [24] $**{}^{\text{*}}$- $J$ is firmly nonexpansive.
PnP-PDS Ono[25] $**{}^{\text{*}}$- -- --
Garcia et al. [26] Constraint $J$ is firmly nonexpansive.
Ours $**{}^{\text{*}}$ $J$ is firmly nonexpansive.
Data-fidelity functions should be proper lower-semicontinuous convex, and their proximal operators should have closed-form expressions or be efficiently computable.
Recently, regularization using Gaussian denoisers has become an active area of research. This line of work was pioneered by a PnP paradigm that was proposed in [19] and later extended, e.g., in [27, 28, 20]. The basic idea of PnP approaches is to replace the proximity operator of the regularization term used in proximal splitting algorithms with an off-the-shelf Gaussian denoiser, such as Block-Matching and 3D Filtering (BM3D) [29], Non-Local Means (NLM) [30], Trainable Nonlinear Reaction Diffusion (TNRD) [31], and Denoising Convolutional Neural Networks (DnCNN)[32]. Numerous studies have shown that the PnP strategy achieves superior performance across a variety of applications, including tomographic imaging [19, 21], biomedical imaging [33], magnetic resonance imaging [34], and remote sensing [35] (see [36] for a recent review).
Following these impressive results, several works explore attempts to plug in more sophisticated denoisers. Zhang et al. proposed a neural network specifically designed for use in PnP methods [37], known as DRUNet. Duff et al. proposed generative regularizers [38], realized by Autoencoders (AE), Variational Autoencoders (VAE)[39], and Generative Adversarial Networks (GAN) [40]. Zhu et al. showed that diffusion models [41] also have promising performance as prior knowledge used in PnP methods [42]. We should mention that despite such extensive research, the theoretical convergence guarantee for PnP methods remains a challenging open question [43, 21, 22, 18]. This is mainly because replacing the proximity operator of the regularization term with a denoiser does not necessarily guarantee the existence of the underlying convex optimization problem.
To explore explicit data-driven regularization, alternative frameworks including regularization by denoising (RED) [44, 45], gradient step denoisers (GSD) [46, 47, 48, 49], and its generalization in Bregman geometry [50] have been proposed. Unlike PnP strategies, they explicitly define a regularization term derived from the output of a denoiser, incorporating it directly into the formulation of the optimization problem. Subsequently, they construct algorithms to solve this problem as a special case of traditional proximal splitting algorithms. Nonetheless, it should be noted that they still rely on conditions that may not be desirable in real-world settings. For example, RED imposes strict assumptions on the denoiser to guarantee its convergence, which are often impractical for deep denoisers [51] (see Section III-C4 for details). The GSD approach introduces a non-convex regularization term, which prevents guarantees of convergence to globally optimal solutions and makes its performance unreliable in practice.
In contrast, within the context of PnP methods, some studies have been conducted to establish their convergence under realistic assumptions, focusing on the firmly nonexpansiveness of denoisers. Pesquet et al. showed theoretically that a subclass of maximally monotone operators (MMO) can be fully represented by the firmly nonexpansive DnCNN with a certain network structure. They also constructed a PnP method based on forward-backward splitting (PnP-FBS) with convergence guarantees for image restoration [17]. As highlighted by the authors, their technique can be extended to any algorithm, as long as its analysis relies on the theory of MMOs, including PDS. Sun et al. proposed a PnP based on ADMM (PnP-ADMM) with the convergence guarantee by plugging in a firmly nonexpansive denoiser [24]. Hertrich et al. [52] developed a method to train averaged CNNs by stochastic gradient descent on the Stiefel manifold.
Regarding the flexibility of PnP algorithms, PnP based on PDS (PnP-PDS) offers the advantage of efficiently handling a wider range of image restoration problems than PnP-FBS and PnP-ADMM, as mentioned above. Ono proposed the first PnP-PDS and demonstrated its practicality, while no theoretical convergence guarantees were provided [25]. On the other hand, Garcia et al. proposed a convergent PnP-PDS with a firmly nonexpansive denoiser for a specific case and achieved successful results in the restoration of images obtained by a photonic lantern [26]. We remark that the denoiser assumption of firmly nonexpansiveness is reasonable in actual applications, as shown in [17]. Nevertheless, the convergence of PnP-PDS in its general form has not yet been fully analyzed. Once a general PnP-PDS with convergence guarantees is established, various image restoration problems can be approached with rich priors of Gaussian denoisers in an efficient and flexible fashion.
I-C Contributions and Paper Organization
Now, the following natural question arises: Can we construct a general PnP-PDS with theoretical convergence guarantees under realistic assumptions on denoisers? In this paper, we answer this question by proposing a general convergent PnP-PDS strategy with a theoretical guarantee that has a wide range of practical applications. The main contributions of our paper are outlined below:
  • We study a general PnP-PDS algorithm that extends the framework proposed in [25] and establish its theoretical convergence conditions, building upon the foundational results presented in [17].
  • We show that the problem solved by our PnP-PDS algorithm is not a standard convex optimization problem but a certain monotone inclusion problem, and provide a mathematical representation of the solution set.
  • Based on the above foundations, we show that our PnP-PDS efficiently solves typical inverse problems formulated with non-quadratic data-fidelity, with convergence guarantees. In particular, it allows the use of an $\ell_{2}$-ball data-fidelity constraint for image restoration under Gaussian noise, offering more flexibility in parameter selection and improving the versatility of the denoiser. Moreover, it effectively handles the generalized Kullback–Leibler (GKL) divergence for image restoration under Poisson noise.
In Table I, we briefly summarize how the proposed method differs from related works (see Section III-C for detailed and comprehensive discussion). The advantages of the proposed convergent PnP-PDS are illustrated by experimental results on a number of typical image restoration tasks.
The remainder of this paper is organized as follows. In Section II, we set up proximal tools, the PDS algorithm, and the firmly nonexpansive denoiser to construct our PnP-PDS. Section III presents our main results. We first establish a convergence guarantee for the general PnP-PDS algorithm and characterize its solutions. Subsequently, we introduce several specific forms of the algorithm designed for image restoration tasks, along with their corresponding solution characterizations. In Section IV, we conduct image restoration experiments, focusing on deblurring and inpainting tasks under two noise conditions: Gaussian noise and Poisson noise, and demonstrate that our PnP-PDS achieves state-of-the-art performance and high stability.
The preliminary version of this paper, without the applications to deblurring and inpainting under Poisson noise, the introduction of a box constraint, and the complete theoretical proof, can be found in the conference proceedings [53].
II Preliminaries
II-A Notations
Let $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ denote the $\ell_{1}$-norm and $\ell_{2}$-norm, respectively. The transpose of a vector is denoted by $(\cdot)^{\top}$. The notation $\left[\cdot\right]_{m}$ indicates the $m$-th element of a vector. The adjoint of a bounded linear operator is denoted by $(\cdot)^{*}$, and the operator norm is represented as $\|\cdot\|_{\mathrm{op}}$.
[…]
For $D\subset\mathcal{H}$, we consider an operator $T:D\to\mathcal{H}$. If $\|T\mathbf{x}_{1}-T\mathbf{x}_{2}\|\leq\|\mathbf{x}_{1}-\mathbf{x}_{2}\|$ holds for every $(\mathbf{x}_{1},\,\mathbf{x}_{2})\in D^{2}$, $T$ is said to be nonexpansive. When $2T-\mathrm{Id}$ is nonexpansive, $T$ is called firmly nonexpansive [54, Proposition 4.4] ($\mathrm{Id}$ denotes an identity operator). If $\kappa T$ is firmly nonexpansive for some $\kappa\in(0,\,\infty)$, then $T$ is $\kappa$-cocoercive.
Let $A:\mathcal{H}\to 2^{\mathcal{H}}$ be a set-valued operator. Its graph is $\mathrm{gra}A=\{(\mathbf{x},\,\mathbf{y})\in\mathcal{H}^{2}:\mathbf{y}\in A\mathbf{x}\}$. If $\left\langle\mathbf{x}_{1}-\mathbf{x}_{2},\,\mathbf{y}_{1}-\mathbf{y}_{2}\right\rangle\geq 0$ holds for every $(\mathbf{x}_{1},\,\mathbf{y}_{1})\in\mathrm{gra}A$ and $(\mathbf{x}_{2},\,\mathbf{y}_{2})\in\mathrm{gra}A$, $A$ is said to be monotone. The operator $A$ is called maximally monotone if there exists no monotone operator $B:\mathcal{H}\to 2^{\mathcal{H}}$ such that $\mathrm{gra}B$ properly contains $\mathrm{gra}A$.
We introduce several specific convex functions and sets that we use in this paper. Let $C$ be a nonempty closed convex set on $\mathcal{H}$. The indicator function of $C$, denoted by $\iota_{C}$, is defined as
$\iota_{C}(\mathbf{x}):=\begin{cases}0,&\mathrm{if}\,\mathbf{x}\in C,\\ \infty,&\mathrm{otherwise}.\end{cases}$ (1)
Let $B_{2,\,\varepsilon}^{\mathbf{v}}$ be the $\mathbf{v}$-centered $\ell_{2}$-norm ball with radius $\varepsilon>0$, which is given by
$B_{2,\,\varepsilon}^{\mathbf{v}}:=\left\{\mathbf{x}\in\mathcal{H}\mid\|\mathbf{x}-\mathbf{v}\|_{2}\leq\varepsilon\right\}.$ (2)
The Generalized Kullback-Leibler (GKL) divergence between $\mathbf{x}\in\mathbb{R}^{N}$ and $\mathbf{v}\in\mathbb{R}^{N}$ is defined as
$\mathrm{GKL}_{\mathbf{v}}(\mathbf{x})$ (3)
$:=\displaystyle\sum_{i=1}^{N}\begin{cases}\eta[\mathbf{x}]_{i}-[\mathbf{v}]_{i}\ln{\eta[\mathbf{x}]_{i}},\,&\text{if $[\mathbf{v}]_{i}>0$ and $[\mathbf{x}]_{i}>0$},\\ \eta[\mathbf{x}]_{i},\,&\text{if $[\mathbf{v}]_{i}=0$ and $[\mathbf{x}]_{i}\geq 0$,}\\ \infty,\,&\text{otherwise,}\end{cases}$ (4)
where $\eta>0$ is a certain scaling parameter, which will be detailed in Section III-B2.
II-B Proximal Tools
For $\mathcal{F}\in\Gamma_{0}(\mathcal{H})$, the proximity operator of an index $\gamma>0$ is defined as follows[55]:
$\mathrm{prox}_{\gamma\mathcal{F}}(\mathbf{x}):=\underset{\mathbf{y}\in\mathcal{H}}{\operatorname{argmin}}\,\mathcal{F}(\mathbf{y})+\dfrac{1}{2\gamma}\|\mathbf{y}-\mathbf{x}\|_{2}^{2}.$ (5)
Let us consider the convex conjugate function of $\mathcal{F}$, defined as follows:
$\mathcal{F}^{*}(\mathbf{x}):=\underset{\mathbf{\mathbf{y}}\in\mathcal{H}}{\operatorname{sup}}\,\left\langle\mathbf{x},\,\mathbf{y}\right\rangle-\mathcal{F}(\mathbf{y}).$ (6)
According to the Moreau’s identity[56, Theorem 3.1 (ii)], $\mathrm{prox}_{\gamma\mathcal{F}^{*}}$ can be computed via $\mathrm{prox}_{\mathcal{F}/\gamma}$ as
$\mathrm{prox}_{\gamma\mathcal{F}^{*}}(\mathbf{x})=\mathbf{x}-\gamma\mathrm{prox}_{\mathcal{F}/\gamma}(\mathbf{x}/\gamma).$ (7)
The proximity operator of $\iota_{C}$ equals to the metric projection onto $C$, i.e.,
$\mathrm{prox}_{\iota_{C}}(\mathbf{x})=P_{C}(\mathbf{x}):=\underset{\mathbf{y}\in C}{\operatorname{argmin}}\,\|\mathbf{y}-\mathbf{x}\|_{2}.$ (8)
In particular, it can be calculated for $C:=B_{2,\,\varepsilon}^{\mathbf{v}}$ as follows:
$\mathrm{prox}_{\iota_{B_{2,\,\varepsilon}^{\mathbf{v}}}}(\mathbf{x})=\begin{cases}\mathbf{x},&\mathrm{if}\,\mathbf{x}\in C,\\ \mathbf{v}+\dfrac{\varepsilon(\mathbf{x}-\mathbf{v})}{\|\mathbf{x}-\mathbf{v}\|_{2}},&\mathrm{otherwise}.\end{cases}$ (9)
The proximity operator of the GKL divergence can be computed as follows:
$\left[\mathrm{prox}_{\gamma\mathrm{GKL}_{\mathbf{v}}}(\mathbf{x})\right]_{i}=\dfrac{1}{2}([\mathbf{x}]_{i}-\gamma\eta+\sqrt{([\mathbf{x}]_{i}-\gamma\eta)^{2}+4\gamma[\mathbf{v}]_{i}}).$ (10)
II-C Primal-Dual Splitting Algorithm
Let us review a PDS algorithm [13],[14], which is a basis of our PnP strategy. Let $\mathcal{X}$ and $\mathcal{Y}$ be two real Hilbert spaces. Suppose that $f\in\Gamma_{0}(\mathcal{X})$ is differentiable with its gradient $\nabla f$ being $\beta$-Lipschitz continuous with $\beta>0$, $g\in\Gamma_{0}(\mathcal{X})$ and $h\in\Gamma_{0}(\mathcal{Y})$ are proximable functions, i.e., the proximity operators have a closed form solution or can be computed efficiently, and $\mathbf{L}:\mathcal{X}\to\mathcal{Y}$ is a bounded linear operator. Consider the following convex optimization problem:
$\underset{\mathbf{x}\in\mathcal{X}}{\operatorname{minimize}}\,f(\mathbf{x})+g(\mathbf{x})+h(\mathbf{L}\mathbf{x}),$ (11)
where the set of solutions to (11) is assumed to be nonempty. Its dual formulation is given as follows [54, Chap. 15]:
$\underset{\mathbf{y}\in\mathcal{Y}}{\operatorname{minimize}}\,(f+g)^{*}(-\mathbf{L}^{*}\mathbf{y})+h^{*}(\mathbf{y}).$ (12)
The PDS algorithm jointly solves both (11) and (12) by the following iterative procedure:
$\begin{split}&\mathbf{x}_{n+1}=\mathrm{prox}_{\gamma_{1}g}\left(\mathbf{x}_{n}-\gamma_{1}(\nabla f(\mathbf{x}_{n})+\mathbf{L}^{*}\mathbf{y}_{n})\right),\\ &\mathbf{y}_{n+1}=\mathrm{prox}_{\gamma_{2}h^{*}}\left(\mathbf{y}_{n}+\gamma_{2}\mathbf{L}(2\mathbf{x}_{n+1}-\mathbf{x}_{n})\right),\end{split}$ (13)
where $\gamma_{1}>0$ and $\gamma_{2}>0$ are parameters that satisfy
$\gamma_{1}\left(\dfrac{\beta}{2}+\gamma_{2}\|\mathbf{L}\|_{\mathrm{op}}^{2}\right)<1.$ (14)
Additionally, we consider the following monotone inclusion problem derived from the optimality conditions of (11) and (12):
$\begin{split}\mathrm{Find}\,&(\hat{\mathbf{x}},\,\hat{\mathbf{y}})\in\,\mathcal{X}\times\mathcal{Y}\\ \mathrm{s.t.}&\begin{pmatrix}\mathbf{0}\\ \mathbf{0}\end{pmatrix}\in\begin{pmatrix}\partial g(\hat{\mathbf{x}})+\mathbf{L}^{*}\hat{\mathbf{y}}+\nabla f(\hat{\mathbf{x}})\\ -\mathbf{L}\hat{\mathbf{x}}+\partial h^{*}(\hat{\mathbf{y}})\end{pmatrix},\end{split}$ (15)
where ($\partial$($\cdot$)∂ ( ⋅ ) denotes the subdifferential of a convex function. Under a certain mild condition, the solutions to (15) coincide with the solutions to (11) and (12) (see [13] and [14] for details).
II-D Firmly Nonexpansive Denoiser[17]
We review the method proposed in [17] for training a firmly nonexpansive denoiser represented by DnCNN, which is essential for the construction of convergent PnP algorithms. Basically, they add a specific penalty during training to ensure the firmly nonexpansiveness of the denoiser.
Let $J:\mathbb{R}^{K}\to\mathbb{R}^{K}$ be a Gaussian denoiser represented by DnCNN. To train a firmly nonexpansive denoiser, we first consider the operator $Q:=2J-\mathrm{Id}$, which is assumed to be differentiable.
By the definition of firmly nonexpansiveness, $J$ is firmly nonexpansive if and only if $Q$ is nonexpansive. Thus, our objective is to ensure the nonexpansiveness of $Q$ rather than the firmly nonexpansiveness of $J$. If $Q$ is differentiable, its nonexpansiveness is equivalent to its Jacobian $\nabla Q$ satisfying the following condition for all $\mathbf{x}\in\mathbb{R}^{K}$:
$\|\nabla Q(\mathbf{x})\|_{\mathrm{sp}}\leq 1,$ (16)
where $\|\cdot\|_{\mathrm{sp}}$ represents the spectral norm.
Based on the above fact, the authors of [17] proposed a penalty to encourage $Q$ to satisfy the condition in (16), and defined a new loss function expressed by the sum of the squared prediction error and the penalty. Specifically, the loss function is calculated for each image as
[…]
Here, $\tau>0$ is a penalization parameter, $\xi>0$ is a parameter that defines a safety margin, $\overline{\mathbf{x}}_{\ell}$ is the $\ell$-th ground truth image in the training dataset, $\mathbf{x}_{\ell}$ is the image corrupted with Gaussian noise, and $\tilde{\mathbf{x}}_{\ell}$ is the image calculated by the following equation:
$\tilde{\mathbf{x}}_{\ell}=\rho\overline{\mathbf{x}}_{\ell}+(1-\rho)J(\mathbf{x}_{\ell}),$ (18)
where $\rho$ is a random value chosen according to a uniform distribution in the range from 00 to $1$.
Note that $\|\nabla Q(\mathbf{\tilde{x}}_{\ell})\|_{\mathrm{sp}}$ can be computed using the power iterative method[57]. For more detailed information on the training method, see [17, Section 3.2].
III Proposed Method
We propose a general convergent PnP-PDS, demonstrate its application to image restoration with specific cases of our PnP-PDS, and provide detailed discussions on its differences from related methods. For clarity, we present a flowchart in Fig 1 that illustrates the relationships between our theoretical results.
Figure 1: A flowchart showing the basic structure of the theoretical discussions in this paper. The dotted and solid lines represent existing and new results, respectively.
III-A General Form of Convergent PnP-PDS and Its Convergence Analysis
We propose a convergent PnP-PDS that substitutes the proximity operator of $g$ in (LABEL:PDSupdate) with a firmly nonexpansive denoiser $J$ represented by DnCNN. The iterative procedures of our PnP-PDS is given by
$\begin{split}&\widetilde{\mathbf{x}}_{n+1}=J\left(\mathbf{x}_{n}-\gamma_{1}(\nabla f(\mathbf{x}_{n})+\mathbf{L}^{*}\mathbf{y}_{n})\right),\\ &\widetilde{\mathbf{y}}_{n+1}=\mathrm{prox}_{\gamma_{2}h^{*}}\left(\mathbf{y}_{n}+\gamma_{2}\mathbf{L}(2\mathbf{x}_{n+1}-\mathbf{x}_{n})\right),\\ &\mathbf{x}_{n+1}=\rho_{n}\widetilde{\mathbf{x}}_{n+1}+(1-\rho_{n})\mathbf{x}_{n},\\ &\mathbf{y}_{n+1}=\rho_{n}\widetilde{\mathbf{y}}_{n+1}+(1-\rho_{n})\mathbf{y}_{n},\end{split}$ (19)
where $\gamma_{1},\,\gamma_{2}$ and $\{\rho_{n}\}$ are the parameters and the definitions of $f$, $h$, and $\mathbf{L}$ are the same as in (11).
The following two propositions present our main theoretical results on the convergence of PnP-PDS in (19). Proposition 3.1 corresponds to the case with the differentiable term $f$, while Proposition 3.2 corresponds to the case without it.
Proposition 3.1.
Suppose that $\beta>0$. Let $J:\mathcal{X}\to\mathcal{X}$ be a firmly nonexpansive denoiser and defined over the full domain. Then, there exists a maximally monotone operator $A_{J}$ such that $J=(A_{J}+\mathrm{Id})^{-1}$. Assume that the solution set of the following monotone inclusion problem:
$\begin{split}\mathrm{Find}\,&(\hat{\mathbf{x}},\,\hat{\mathbf{y}})\in\,\mathcal{X}\times\mathcal{Y}\\ &\mathrm{s.t.}\begin{pmatrix}\mathbf{0}\\ \mathbf{0}\end{pmatrix}\in\begin{pmatrix}\gamma_{1}^{-1}A_{J}(\hat{\mathbf{x}})+\mathbf{L}^{*}\hat{\mathbf{y}}+\nabla f(\hat{\mathbf{x}})\\ -\mathbf{L}\hat{\mathbf{x}}+\partial h^{*}(\hat{\mathbf{y}})\end{pmatrix}\end{split}$ (20)
is nonempty, and suppose that the following conditions hold:
  1. $(i)$
    $\dfrac{1}{\gamma_{1}}-\gamma_{2}\|\mathbf{L}\|_{\mathrm{op}}^{2}>\dfrac{\beta}{2}$,
  2. $(ii)$
    $\rho_{n}\in\left(0,\,\delta\right)$,
  3. $(iii)$
    $\sum_{n\in\mathbb{N}}\rho_{n}(\delta-\rho_{n})=\infty$,
where $\gamma_{1}>0$, $\gamma_{2}>0$, and $\delta$ is defined as follows:
$\delta:=2-\dfrac{\beta}{2}\left(\dfrac{1}{\gamma_{1}}-\gamma_{2}\|\mathbf{L}\|_{\mathrm{op}}^{2}\right)^{-1}.$ (21)
Then, the sequence $\{\mathbf{x}_{n},\,\mathbf{y}_{n}\}_{n\in\mathbb{N}}$ generated by (19) weakly converges to a solution to (20).
Proposition 3.2.
Suppose that $f=0$. Let $J:\mathcal{X}\to\mathcal{X}$ be a firmly nonexpansive denoiser and defined over the full domain. Then, there exists a maximally monotone operator $A_{J}$ such that $J=(A_{J}+\mathrm{Id})^{-1}$. Assume that the solution set of (20) is nonempty, and the following conditions hold:
  1. $(i)$
    $\dfrac{1}{\gamma_{1}}-\gamma_{2}\|\mathbf{L}\|_{\mathrm{op}}^{2}>0$,
  2. $(ii)$
[…]
  3. $(iii)$
    $\sum_{n\in\mathbb{N}}\rho_{n}(2-\rho_{n})=\infty$,
where $\gamma_{1}>0$ and $\gamma_{2}>0$. Then, the sequence $\{\mathbf{x}_{n},\,\mathbf{y}_{n}\}_{n\in\mathbb{N}}$ generated by (19) weakly converges to a solution to (20).
Proof Sketch of Propositions 3.1 and 3.2.
Recall that the set-valued operator $A_{J}=J^{-1}-\mathrm{Id}$ is maximally monotone if and only if $J$ is firmly nonexpansive and has the full domain [54, Proposition 23.8 (iii)]. Hence, $\gamma_{1}^{-1}A_{J}$ is also maximally monotone [54, Proposition 20.22]. Therefore, Proposition 3.1 and Proposition 3.2 are justified by [13, Theorem 3.1] and [13, Theorem 3.2], respectively. These proofs build upon and directly apply the standard convergence analysis of the PDS algorithm. For completeness and ease of reference, we provide the detailed proofs of these propositions in the Appendix. ∎
Remark 1 (Connection to the convex optimization problem in (11)).
The monotone inclusion problem in (20) has the same form as (15) with $\partial g=\gamma_{1}^{-1}A_{J}$. The classical monotone operator theory states that there exists $g\in\Gamma_{0}(\mathbb{R}^{N})$ such that $\partial g=\gamma_{1}^{-1}A_{J}$ if and only if $\gamma_{1}^{-1}A_{J}$ is a maximal cyclically monotone operator, a specific type of maximally monotone operator [54, Theorem 22.18]. However, the operator $\gamma_{1}^{-1}A_{J}$, where $J=(A_{J}+\mathrm{Id})^{-1}$, is not necessarily maximal cyclically monotone only because $J$ is firmly nonexpansive. Thus, the existence of $g\in\Gamma_{0}(\mathbb{R}^{N})$ cannot always be guaranteed. Furthermore, it should be noted that $A_{J}$ is scaled by $\gamma_{1}^{-1}$ in the solution set of the monotone inclusion problem (20). This implies that the value of $\gamma_{1}$ affects the solution set, unlike the standard PDS algorithm.
III-B Applications to Image Restoration
Now we move on to the applications of our general PnP-PDS to image restoration. First, we derive a special case of our PnP algorithm in (19), which is set up for typical image restoration problems. Then, we demonstrate its application to specific image restoration tasks with two types of noise contamination: Gaussian and Poisson. In the following, $\mathbf{u}\in\mathbb{R}^{K}$ and $\mathbf{v}\in\mathbb{R}^{K}$ denote original and observed images, respectively. Let $\mathbf{\Phi}:\mathbb{R}^{K}\to\mathbb{R}^{K}$ be an operator representing some observation process, which may include blurring or random sampling. We assume that the pixel values of each image are normalized to the range of 0 to 1.
First, we introduce an observation model that covers many image restoration scenarios as follows:
$\mathbf{v}=\mathcal{N}(\mathbf{\Phi}\mathbf{u}),$ (22)
where $\mathcal{N}:\mathbb{R}^{K}\to\mathbb{R}^{K}$ represents noise contamination (not necessarily additive). Image restoration under (22) can be formulated as the following optimization problem:
$\begin{split}\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\mathcal{D}(\mathbf{\Phi}\mathbf{\hat{u}})+\mathcal{R}_{J}(\mathbf{\hat{u}})\quad\mathrm{s.t.}\,\mathbf{\hat{u}}\in C,\end{split}$ (23)
or equivalently,
$\begin{split}\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\mathcal{D}(\mathbf{\Phi}\mathbf{\hat{u}})+\mathcal{R}_{J}(\mathbf{\hat{u}})+\iota_{C}(\mathbf{\hat{u}}),\end{split}$ (24)
where $\mathcal{D}\in\Gamma_{0}(\mathbb{R}^{K})$ is a proximable data-fidelity term derived from the noise model $\mathcal{N}$, $\mathcal{R}_{J}$ is a proximable regularization term, $C$ is a nonempty closed convex set on $\mathbb{R}^{K}$ to which $\hat{\mathbf{v}}$ is assumed to belong, and $\iota_{C}$ is the indicator function of $C$. We also assume that $\mathcal{R}_{J}$ models the prior distribution associated with the Gaussian denoiser $J$ with $\mathrm{prox}_{\mathcal{R}_{J}}=J$, since $\mathrm{prox}_{\mathcal{R}_{J}}$ corresponds to the MAP estimate under additive white Gaussian noise [19]. Note that $\mathcal{R}_{J}$ is introduced for convenience as an explicit regularization term, but the existence of such a function is not guaranteed (see Remark 1).
Then, we can derive a convergent PnP-PDS algorithm for this type of image restoration as a special case of our general PnP-PDS in (19). The algorithm and its convergence property are summarized in the following corollary.
Corollary 1.
Let $J:\mathbb{R}^{K}\to\mathbb{R}^{K}$ be a firmly nonexpansive denoiser and defined over the full domain. Then, there exists a maximally monotone operator $A_{J}$ such that $J=(A_{J}+\mathrm{Id})^{-1}$. Assume that the solution set of the following monotone inclusion problem:
$\begin{split}\mathrm{Find}\,&(\hat{\mathbf{x}},\,\hat{\mathbf{y}}^{(1)},\,\hat{\mathbf{y}}^{(2)})\in\,\mathbb{R}^{3K}\\ &\mathrm{s.t.}\begin{pmatrix}\mathbf{0}\\ \mathbf{0}\\ \mathbf{0}\end{pmatrix}\in\begin{pmatrix}\gamma_{1}^{-1}A_{J}(\hat{\mathbf{x}})+\mathbf{\Phi}^{*}\hat{\mathbf{y}}^{(1)}+\hat{\mathbf{y}}^{(2)}\\ -\mathbf{\Phi}\hat{\mathbf{x}}+\partial\mathcal{D}^{*}\left(\hat{\mathbf{y}}^{(1)}\right)\\ -\hat{\mathbf{x}}+\partial\iota^{*}_{C}\left(\hat{\mathbf{y}}^{(2)}\right)\end{pmatrix}\end{split}$ (25)
is nonempty, and suppose that the following conditions hold for $\gamma_{1}>0$ and $\gamma_{2}>0$:
$\dfrac{1}{\gamma_{1}}-\gamma_{2}\left(\|\mathbf{\Phi}\|_{\mathrm{op}}^{2}+1\right)>0,$ (26)
and $\{\rho_{n}\}$ satisfies the condition in $(ii)$ and $(iii)$ of Proposition 3.2. Then, the sequence $\{\mathbf{x}_{n},\,\mathbf{y}^{(1)}_{n},\,\mathbf{y}^{(2)}_{n}\}_{n\in\mathbb{N}}$ generated by the following algorithm:
$\begin{split}&\widetilde{\mathbf{x}}_{n+1}=J\left(\mathbf{x}_{n}-\gamma_{1}\left(\mathbf{\Phi}^{*}\mathbf{y}_{n}^{(1)}+\mathbf{y}_{n}^{(2)}\right)\right),\\ &\widetilde{\mathbf{y}}_{n+1}^{(1)}=\mathrm{prox}_{\gamma_{2}\mathcal{D}^{*}}\left(\mathbf{y}_{n}^{(1)}+\gamma_{2}\mathbf{\Phi}(2\mathbf{x}_{n+1}-\mathbf{x}_{n})\right),\\ &\widetilde{\mathbf{y}}_{n+1}^{(2)}=\mathrm{prox}_{\gamma_{2}\iota_{C}^{*}}\left(\mathbf{y}_{n}^{(2)}+\gamma_{2}(2\mathbf{x}_{n+1}-\mathbf{x}_{n})\right),\\ &\mathbf{x}_{n+1}=\rho_{n}\widetilde{\mathbf{x}}_{n+1}+(1-\rho_{n})\mathbf{x}_{n},\\ &\mathbf{y}_{n+1}^{(1)}=\rho_{n}\widetilde{\mathbf{y}}_{n+1}^{(1)}+(1-\rho_{n})\mathbf{y}_{n}^{(1)},\\ &\mathbf{y}_{n+1}^{(2)}=\rho_{n}\widetilde{\mathbf{y}}_{n+1}^{(2)}+(1-\rho_{n})\mathbf{y}_{n}^{(2)},\end{split}$ (27)
converges to a solution to (25).
This is a direct consequence of Proposition 3.2. Let us define $f(\mathbf{x}),\,h\left(\mathbf{y}^{(1)},\,\mathbf{y}^{(2)}\right)$ in (20) as follows:
$f\left(\mathbf{x}\right)$ $:=0,$ (28)
$h(\mathbf{y}^{(1)},\,\mathbf{y}^{(2)})$ $:=\mathcal{D}(\mathbf{y}^{(1)})+\iota_{C}(\mathbf{y}^{(2)}),$ (29)
$\mathbf{L}:$ $=(\mathbf{\Phi},\,\mathbf{I}),$ (30)
where $\mathbf{I}$ represents an identity matrix. Then, the monotone inclusion problem in (20) is reduced to (25) [54, Proposition 13.30, Proposition 16.9, Corollary 16.48], and we obtain Algorithm (27) from (19). The condition regarding $\gamma_{1}$ and $\gamma_{2}$ is obtained from the assumption in $(i)$ of Proposition 3.2 and the following fact [58, Theorem 1]:
$\|\mathbf{L}\|_{\mathrm{op}}^{2}\leq\|\mathbf{\Phi}\|_{\mathrm{op}}^{2}+\|\mathbf{I}\|_{\mathrm{op}}^{2}=\|\mathbf{\Phi}\|_{\mathrm{op}}^{2}+1.$ (31)
III-B1 Image Restoration under Gaussian Noise
Consider the observation model in (22) with $\mathcal{N}(\mathbf{x}):=\mathbf{x}+\mathbf{n}$, where $\mathbf{n}\in\mathbb{R}^{K}$ is additive white Gaussian noise. Image restoration under this observation model can be formulated as the following optimization problem:
[…]
where $\varepsilon>0$ is the radius of the $\ell_{2}$ data-fidelity constraint, which can be determined by the standard deviation of $\mathbf{n}$.
To apply the algorithm in (27), we define $\mathcal{D}(\cdot)$ and $C$ in (23) as follows:
$\mathcal{D}(\cdot)$ $:=\iota_{B_{2,\,\varepsilon}^{\mathbf{v}}}(\cdot),$ (33)
[…]
where $B_{2,\,\varepsilon}^{\mathbf{v}}$ is a $\mathbf{v}$-centered $\ell_{2}$-norm ball with the radius $\varepsilon$. Then, the problem in (32) can be expressed as the form of (23), allowing the application of (27). If we set $\rho_{n}=1$ for all $n\in\mathbb{N}$, we obtain Algorithm 1.
Remark 2 (Constrained formulation of data-fidelity).
In (32), data-fidelity is expressed as a hard constraint using the $\ell_{2}$-norm ball. The benefits of expressing data-fidelity or regularization terms as hard constraints have been illustrated in the literature [59, 60, 61, 62, 63]. We will discuss such advantages in the specific case of (32) for PnP methods in Section III-C1.
Algorithm 1 Image restoration under Gaussian noise by PnP-PDS.
1:$\mathbf{u}_{0},\,\mathbf{w}_{0}^{(1)},\,\mathbf{w}_{0}^{(2)},\,\gamma_{1}>0,\,\gamma_{2}>0,\,\varepsilon>0$
2:while a stopping criterion is not satisfied do
3:     $\mathbf{u}_{n+1}\leftarrow J\left(\mathbf{u}_{n}-\gamma_{1}\left(\mathbf{\Phi}^{*}\mathbf{w}_{n}^{(1)}+\mathbf{w}_{n}^{(2)}\right)\right)$;
4:     $\mathbf{w}_{n}^{(1)}\leftarrow\mathbf{w}_{n}^{(1)}+\gamma_{2}\mathbf{\Phi}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
5:     $\mathbf{w}_{n}^{(2)}\leftarrow\mathbf{w}_{n}^{(2)}+\gamma_{2}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
6:     $\mathbf{w}_{n+1}^{(1)}\leftarrow\mathbf{w}_{n}^{(1)}-\gamma_{2}P_{B^{\mathbf{v}}_{2,\,\varepsilon}}(\frac{1}{\gamma_{2}}\mathbf{w}_{n}^{(1)})$;
[…]
8:end while
9:$\mathbf{u}_{n}$
III-B2 Image Restoration under Poisson Noise
Image restoration under Poisson noise has been studied across various domains, including medical imaging, astronomical imaging, and remote sensing. The observation process considering Poisson noise contamination can be modeled by (22) when $\mathcal{N}(\mathbf{x}):=\mathcal{P}_{\eta}(\mathbf{x})$, where $\mathcal{P}_{\eta}$ represents the corruption by Poisson noise with the scaling coefficient $\eta$.
In this observation model, the data-fedelity is expressed as the GKL divergence (refer to [64]), and image restoration can be reduced to the following optimization problem:
[…]
where $\lambda>0$ is a parameter that determines the balance between data-fidelity and regularization (see (4) for the definition of $\mathrm{GKL}$).
Let us define $\mathcal{D}(\cdot)$ and $C$ in (23) as follows:
$\mathcal{D}(\cdot)$ $:=\lambda\mathrm{GKL}_{\mathbf{v}}(\cdot),$ (36)
[…]
Then, the problem in (35) is reformulated as (23), and the iterative procedure in (27) becomes applicable. Setting $\rho_{n}=1$ for all $n\in\mathbb{N}$, the resulting algorithm is given as Algorithm 2.
Algorithm 2 Image restoration under Poisson noise by PnP-PDS.
1:$\mathbf{u}_{0},\,\mathbf{w}_{0}^{(1)},\,\mathbf{w}_{0}^{(2)},\,\gamma_{1}>0,\,\gamma_{2}>0,\,\lambda>0$
2:while a stopping criterion is not satisfied do
3:     $\mathbf{u}_{n+1}\leftarrow J\left(\mathbf{u}_{n}-\gamma_{1}\left(\mathbf{\Phi}^{*}\mathbf{w}_{n}^{(1)}+\mathbf{w}_{n}^{(2)}\right)\right)$;
4:     $\mathbf{w}_{n}^{(1)}\leftarrow\mathbf{w}_{n}^{(1)}+\gamma_{2}\mathbf{\Phi}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
5:     $\mathbf{w}_{n}^{(2)}\leftarrow\mathbf{w}_{n}^{(2)}+\gamma_{2}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
6:     $\mathbf{w}_{n+1}^{(1)}\leftarrow\mathbf{w}_{n}^{(1)}-\gamma_{2}\mathrm{prox}_{\frac{\lambda}{{\gamma_{2}}}\mathrm{GKL}}(\frac{1}{\gamma_{2}}\mathbf{w}_{n}^{(1)})$;
[…]
8:end while
9:$\mathbf{u}_{n}$
III-C Comparison with Prior Works
III-C1 Convergent PnP-FBS [17]
Pesquet et al. proposed a PnP-FBS algorithm using DnCNN as a convergent PnP approach[17]. This method can be applied to image restoration under Gaussian noise, addressing the following optimization problem:
$\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\frac{\lambda}{2}\|\mathbf{\Phi\hat{u}}-\mathbf{v}\|_{2}^{2}+\mathcal{R}_{J}(\mathbf{\hat{u}}),$ (38)
where the first term is an $\ell_{2}$ data-fidelity term, and $\lambda>0$ is a balancing parameter.
This convergent PnP-FBS is a landmark method that first ensures the convergence of PnP-FBS with the characterization of the solution set under realistic assumptions. At the same time, it should be noted that the problem in (32), where the data-fidelity term is represented as a hard constraint, cannot be solved by PnP-FBS.
Consequently, the proposed PnP-PDS (Algorithm 1) inherits the versatility of a once-trained denoiser in this setting, compared to PnP-FBS. This advantage arises from its high flexibility in parameter selection. In the case of the constrained formulation in (32), a natural choice for $\varepsilon$ is given as follows [60]:
$\varepsilon=\sigma\sqrt{K},$ (39)
where $\sigma$ denotes the standard deviation of the Gaussian noise in the observed images. We remark that the convergence of the proposed PnP-PDS is guaranteed regardless of the choice of $\varepsilon$.
In contrast, PnP-FBS limits the flexibility of the denoiser in certain cases, due to constraints on the parameter to ensure convergence. Assume that the denoiser is trained with a noise level of $\sigma_{J}$, and that $\mathbf{\Phi}$ represents a circular convolution with kernel $h$. In the additive formulation in (38), the authors suggest the following value for $\lambda$ [17]:
$\lambda_{\mathrm{opt}}:=\frac{\sigma_{J}}{2\sigma\|h\|}.$ (40)
However, $\lambda$ must also satisfy the condition to ensure the convergence of PnP-FBS:
$\lambda<\frac{2}{\|\boldsymbol{\Phi}\|_{\mathrm{op}}^{2}}.$ (41)
Hence, if $\lambda_{\mathrm{opt}}$ does not satisfy the constraint in (41), the convergence of PnP-FBS cannot be ensured with its optimal parameter. This limitation was observed in our experiments (see Section IV-B1).
Another limitation of PnP-FBS is its relatively narrow applicability to problem formulations. Specifically, it does not handle additional constraints explicitly, such as the box constraint in (32), and cannot be applied to the Poisson noise setting in (35), which is formulated with the generalized Kullback–Leibler (GKL) divergence.
III-C2 Convergent PnP-ADMM [24, 23]
Several works have proposed PnP algorithms based on ADMM (PnP-ADMM) and established their convergence guarantees. Nair et al. have proposed a convergent PnP-ADMM and provided extensive analysis on its fixed-point [23]. Sun et al. have proposed a convergent PnP-ADMM algorithm that handles multiple data-fidelity terms, making it practical for large scale datasets [24]. In both studies, the convergence properties are ensured under the firmly nonexpansiveness of the denoiser. PnP-ADMM can address a wider range of tasks than PnP-FBS, including image restoration with Poisson noise.
The main advantage of the proposed PnP-PDS over PnP-ADMM is the elimination of inner iterations. To solve a subproblem in PnP-ADMM algorithms, it is often necessary to perform inner iterations, which results in increasing computational cost and unstable numerical convergence. On the other hand, the proposed PnP-PDS does not need to perform them thanks to the fact that the computations associated with the linear operator $L$ are fully decoupled from the proximity operator of $h^{*}$ in the algorithm.
III-C3 Convergent PnP-PDS[26]
Garcia et al. proposed a convergent PDS-based PnP algorithm for restoration of images obtained by a photonic lantern [26], published contemporaneously with our conference paper [53]. They consider the observation process $\mathbf{v}=\mathbf{\Phi}\mathbf{u}+\mathbf{n}$, where $\mathbf{v}$ is an observed image, $\mathbf{u}$ is an original image, $\mathbf{\Phi}$ is a linear degradation process, and $\mathbf{n}\in\mathbb{R}^{K}$ is an additive noise with an unknown distribution, which is assumed to satisfy $\|\mathbf{n}\|_{2}<\varepsilon$ for $\varepsilon>0$. Based on this observation process, they address the following optimization problem by a convergent PnP-PDS:
$\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\mathcal{R}_{J}(\mathbf{\hat{u}})\quad\mathrm{s.t.}\,\|\mathbf{\Phi}\mathbf{\hat{u}}-\mathbf{v}\|_{2}\leq\varepsilon.$ (42)
We show the relationship between this algorithm and our PnP-PDS algorithm in Fig 1. In essence, the algorithm in [26] is a special case of our Algorithm 1 without the box constraint.
We note that introducing a box constraint ensures that the solution to the optimization (or monotone inclusion) problem remains within a certain convex set, which in turn promotes more stable behavior of PnP iterations for some experimental settings. The experimental results in Section IV illustrate this effect.
III-C4 Regularization by Denoising (RED)[44, 51]
Romano et al. proposed a framework called RED, which defines an explicit objective function by incorporating the output of a denoiser. For image restoration under Gaussian noise and image restoration under Poisson noise, the optimization problems with RED are given respectively as
$\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\lambda\|\mathbf{\Phi}\mathbf{\hat{u}}-\mathbf{v}\|^{2}_{2}+\mathbf{\hat{u}}^{\top}(\mathbf{\hat{u}}-\mathfrak{G}(\mathbf{\hat{u}})),$ (43)
$\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\lambda\mathrm{GKL}_{\mathbf{v}}(\hat{\mathbf{u}})+\mathbf{\hat{u}}^{\top}(\mathbf{\hat{u}}-\mathfrak{G}(\mathbf{\hat{u}})),$ (44)
where $\mathfrak{G}$ represents a Gaussian denoiser. In RED, the optimization algorithms are constructed by computing the gradient of the second terms in (43) and (44) using the following expression:
$\nabla(\mathbf{\hat{u}}^{\top}(\mathbf{\hat{u}}-\mathfrak{G}(\mathbf{\hat{u}})))=\hat{\mathbf{u}}-\mathfrak{G}(\mathbf{\hat{u}}).$ (45)
For differentiable $\mathfrak{G}$, this equation holds if and only if $\mathfrak{G}$ is locally homogeneous and has a symmetric Jacobian [44, 51]. For non-differentiable $\mathfrak{G}$, a more recent study [45] demonstrates that (45) still holds under the assumption that $\mathrm{Id}-\mathfrak{G}$ is maximally cyclically monotone, or that $\mathfrak{G}$ is cyclically firmly nonexpansive. Furthermore, if $\mathfrak{G}$ is also nonexpansive in addition to either of these conditions, we can establish a convergent algorithm based on gradient computations, such as gradient descent or FBS.
The main advantage of using RED is the simplicity resulting from the existence of a clear convex objective function. In contrast, PnP algorithms including our PnP-PDS do not necessarily have a convex objective function. Instead, they solve a monotone inclusion problem expressed as (20).
However, RED also has a considerable difficulty: the assumptions for the denoiser. First, local homogeneity is an unrealistic property, especially for deep denoisers. In fact, the gradients of the regularization terms computed in the RED algorithms often contain substantial computational errors due to the lack of local homogeneity in the denoisers [51]. Second, cyclically firmly nonexpansiveness [45] is even stronger assumption than the firmly nonexpansiveness assumed in our work. These challenges make it difficult to ensure the convergence of RED algorithms when we employ state-of-the-art denoisers, including DnCNN.
IV Experiments
[…]
We performed two types of experiments to demonstrate the stability, performance, and versatility of the proposed method. The first experiment involves deblurring/inpainting under Gaussian noise, and the second experiment involves deblurring/inpainting under Poisson noise. The purpose of these experiments is to confirm the following two facts:
  • The proposed method operates stably and converges in all experimental settings.
  • In comparison to other state-of-the-art methods, the proposed method demonstrates higher restoration performance thanks to its stability.
IV-A Experimental Setup
In the two experiments, we considered two types of observation operators, one for blur and the other for random sampling. For the blur case, the operator $\mathbf{\Phi}$ was defined as a circular convolution with a $19\times 19$ kernel size [66, kernel 1]. This kernel was normalized so that $\|\mathbf{\Phi}\|_{\mathrm{op}}$ was equal to $1$. In the case of random sampling, pixels of the image were randomly selected and masked. The random sampling rate was set to $80\%$ for all cases. The operator norm also satisfies $\|\mathbf{\Phi}\|_{\mathrm{op}}=1$.
We evaluated the restoration performance by Peak Signal to Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM). To investigate stability, the update rate $c_{n}$ for $\0n\in\mathbb{N}\backslash\{0\}$ was defined by $c_{n}:={\|\mathbf{u}_{n}-\mathbf{u}_{n-1}\|_{2}}/{\|\mathbf{u}_{n-1}\|_{2}},$ where $\mathbf{u}_{n}$ is the intermediate solution at the $n$-th iteration by each algorithm.
We used color images for Gaussian noise case and grayscale images for Poisson noise case. The color images were randomly sampled from ImageNet [65], consisting of $7$ images cropped to $128\times 128$ pixels. The grayscale images were selected from 3 widely used image sets [17], all with a size of $256\times 256$ pixels.
Method Algorithm Regularization
PnP-FBS [17] FBS FNE* DnCNN
PnP-ADMM [24] ADMM FNE* DnCNN
PnP-PDS (Unstable) [25] PDS DnCNN (not FNE*)
PnP-PDS (w/o a box const.) [26] PDS FNE* DnCNN
TV [61] PDS TV
RED (Gaussian) [44] SD FNE* DnCNN
RED (Poisson) [44] PDS FNE* DnCNN
DRUNet [37] HQS DRUNet
BPG [50] BPG Bregman Score Denoiser
Proposed PDS FNE* DnCNN
*FNE: firmly nonexpansive   SD: steepest descent
HQS: half-quadratic splitting
TABLE II: Summary of Compared Methods, Their Optimization Algorithms, and Regularization.
In each experiment, we selected the appropriate methods from PnP-FBS [17], PnP-ADMM [24], PnP-PDS (Unstable) [25], PnP-PDS (w/o a box const.) [26], TV [61], RED [44], DRUNet [37], and Bregman Proximal Gradient (BPG) [50], and compared them with the proposed method. The optimization algorithms and regularization for each method are summarized in Table II. Details of PnP-FBS, PnP-ADMM, and RED are given in Section III-C. For BPG, we followed the stopping criterion in the implementation provided by the authors [50]. For all other methods, the number of iterations was fixed at 1200 for the blur case and 3000 for the random sampling case.
For fair comparison, unless otherwise specified, the same DnCNN distributed by Pesquet et al. [17] was used for regularization in PnP-FBS, PnP-ADMM, PnP-PDS (w/o a box const.), RED, and the proposed method. This denoiser was a firmly nonexpansive DnCNN trained on Gaussian noise with a standard deviation of $0.01$. See [17] for details of the training.
Throughout our experiments, the parameters $\gamma_{1}$ and $\gamma_{2}$ in Algorithm 1 and Algorithm 2 were set to $0.5$ and $0.99$, respectively, satisfying the condition in Proposition 3.2. All of our experiments were conducted on a Windows 11 PC, equipped with an Intel Core i9 3.70GHz processor and 32.0GB of RAM.
IV-B Deblurring/Inpainting under Gaussian Noise
We performed experiments on deblurring/inpainting under Gaussian noise. Here, we investigate the versatility of a once-trained denoiser in the proposed method, and compare its performance with existing methods.
IV-B1 Versatility of a Once-Trained Denoiser
The theoretical convergence guarantees in this work build upon PnP-FBS. Hence, we experimentally confirm that the versatility of the denoiser is enhanced by extending PnP-FBS to PnP-PDS.
[…]
Fig. 2 shows the variation in restoration performance across different $\alpha$ and $\zeta$ values. While both methods exhibit comparable peak performance at $\sigma=0.02$ and $\sigma=0.04$, PnP-FBS diverges at lower noise levels (i.e., $\sigma=0.0025$, $0.005$, and $0.01$), as discussed in Section III-C1. For example, in the case of $\sigma=0.01$, we obtain $\lambda_{\mathrm{opt}}=2.22$ from (40) since $0.00452\sigma\|h\|=$ 0.0045 in this situation. In addition, the convergence condition (41) requires $\lambda<2$ since $\|\mathbf{\Phi}\|_{\mathrm{op}}^{2}=1$, which implies $\zeta<0.9$. In fact, stable performance is not observed for $\zeta\geq 0.9$ when $\sigma=0.01$.
In contrast, the proposed method exhibits reasonable performance across all noise levels. This consistency demonstrates the high versatility of the denoiser trained on the specific noise level (see also the discussion in Section III-C1). We also observed that the sensitivity of performance with respect to $\alpha$ tended to increase with higher noise levels, and that the optimal value of $\alpha$ became closer to $1$ as the noise level increased.
IV-B2 Restoration Performance
Noise level $\sigma$ $0.0025$ $0.005$ $0.01$ $0.02$ $0.04$
Deblurring $0.820$ $0.860$ $0.920$ $0.960$ $1.000$
Inpainting $0.900$ $0.820$ $0.820$ $1.000$ $1.000$
TABLE III: The Values of $\alpha$ Used in Each Setting for the Proposed Method.
Deblurring Inpainting
Noise level $\sigma$ $0.0025$ $0.005$ $0.01$ $0.02$ $0.04$ $0.0025$ $0.005$ $0.01$ $0.02$ $0.04$
PnP-FBS [17] $31.28$ $31.16$ $30.59$ $28.19$ $23.95$-- - --- - --- - --- - --- -
PnP-PDS (Unstable) [25] $23.79$ $23.67$ $23.78$ $24.60$ $23.35$ $33.65$ $33.36$ $32.79$ $31.51$ $29.58$
PnP-PDS (w/o a box const.) [26] $\underline{37.23}$ $\underline{34.03}$ $\underline{31.00}$ $28.24$ $\mathbf{26.00}$-- - --- - --- - --- - --- -
TV [61] $31.97$ $30.40$ $28.78$ $27.10$ $25.38$ $\underline{34.19}$ $\underline{33.90}$ $\underline{33.22}$ $\underline{31.90}$ $\underline{29.85}$
RED [44] $34.11$ $33.38$ $30.89$ $\mathbf{28.30}$ $24.87$-- - --- - --- - --- - --- -
DRUNet [37] $32.60$ $31.80$ $28.54$ $-8.04$ $-9.41$ $33.21$ $32.68$ $31.36$ $26.71$ $-15.9$
PnP-PDS (Proposed) $\mathbf{37.24}$ $\mathbf{34.04}$ $\mathbf{31.01}$ $\underline{28.26}$ $\mathbf{26.00}$ $\mathbf{35.61}$ $\mathbf{35.25}$ $\mathbf{34.43}$ $\mathbf{33.04}$ $\mathbf{31.00}$
TABLE IV: The Average PSNR [dB] Values for Deblurring and Inpainting on the 7 Images From ImageNet the Best Value is Highlighted in Bold and the Second Best Value is Underlined.
Figure 3: The evolution of $\|\nabla Q(\mathbf{x}_{n})\|_{\mathrm{sp}}^{2}$ of PnP-PDS (w/o a box constraint), PnP-FBS, RED, and the proposed method for inpainting at $\sigma=0.01$. The values in each plot are computed as the mean over 7 images.
Figure 4: The evolution of PSNR [dB] and $c_{n}$ of PnP-PDS (Unstable) and the proposed method over iterations for deblurring at $\sigma=0.01$. Each plot in the graph corresponds to each image extracted from ImageNet [65].
(a) (PSNR, SSIM)
Ground truth
(b) (21.29, 0.6381)21.290.6381(21.29,\,0.6381)( 21.29 , 0.6381 )
Observation
(c) (33.47, 0.9273)33.470.9273(33.47,\,0.9273)( 33.47 , 0.9273 )
PnP-FBS
(d) (27.55, 0.8465)27.550.8465(27.55,\,0.8465)( 27.55 , 0.8465 )
PnP-PDS
(e) $(\underline{33.93},\,\underline{0.9313})$
PnP-PDS
(w/o a box const.)
(f) (30.53, 0.8875)30.530.8875(30.53,\,0.8875)( 30.53 , 0.8875 )
TV
(g) (33.75, 0.9261)33.750.9261(33.75,\,0.9261)( 33.75 , 0.9261 )
RED
(h) (33.75, 0.9261)33.750.9261(33.75,\,0.9261)( 33.75 , 0.9261 )
DRUNet
(i) (33.97, $(\mathbf{33.97},\,\mathbf{0.9317})$
Proposed
Figure 5: Visual results of image restoration on ImageNet [65] for deblurring at $\sigma=0.01$. The best results are indicated in bold, and the second-best results are underlined.
We compare the performance of the proposed method with other methods, namely PnP-FBS[17], PnP-PDS (Unstable)[25], PnP-PDS (w/o a box const.) [26], TV[61], RED[44], and DRUNet [37]. For constrained formulations such as PnP-PDS (Unstable), PnP-PDS (w/o a box const.), TV, and the proposed method, we set $\varepsilon=\alpha\sigma\sqrt{K}$ in the same way as in the previous section. For each of these methods, we performed a linear search for $\alpha$ to evaluate the restoration performance. In particular, we provide the values of $\alpha$ for the proposed method in Table III. For existing methods, we adopted the recommended values if parameter settings were suggested by the authors. Otherwise, we performed a linear search over the balancing parameter.
Table IV shows the average PSNR values over seven images, obtained with the optimal parameters for each method. The proposed method demonstrates superior performance compared to PnP-FBS, even though both methods employ the same denoiser for regularization. This is attributed to the difference in the versatility of the denoiser, as mentioned in Section IV-B1.
Furthermore, in the case of inpainting by PnP-FBS, PnP-PDS (w/o a box const.), and RED, the algorithms do not converge and PSNR values are not obtained. To discuss this point, let us recall that the denoiser was trained with the penalty in (17), which was designed to promote nonexpansiveness at low-noise images sampled according to (18). During the PnP iterations, however, the inputs to the denoiser can be much more noisy than those images. In such cases, the denoiser may not maintain its firmly nonexpansiveness, which can further amplify instability through successive iterations. For a numerical verification, we present the evolution of $\|\nabla Q(\mathbf{x}_{n})\|_{\mathrm{sp}}^{2}$ at each iteration in Fig. 3. Although PnP-PDS (without a box constraint), PnP-FBS, RED, and the proposed method employ the same denoiser, the values of $\|\nabla Q(\mathbf{x}_{n})\|_{\mathrm{sp}}^{2}$ in PnP-PDS (without a box constraint), PnP-FBS, and RED exhibit unstable behavior and tend to diverge. In contrast, the proposed method consistently maintains values of $\|\nabla Q(\mathbf{x}_{n})\|_{\mathrm{sp}}^{2}$ that satisfy the condition given in (16). While these observations are numerical, they suggest that the denoiser behaves as a firmly nonexpansive operator during the PnP iterations, thanks to the box constraint.
To examine how firmly nonexpansiveness of the denoiser affects stability, we compare PnP-PDS (Unstable) with the proposed method. Fig. 4 shows the evolution of PSNR and the update rate $c_{n}$ at each iteration for the deblurring task with $\sigma=0.01$. For the proposed method, the value of $c_{n}$ decreases to approximately $10^{-5}$ and shows a convergent evolution. In contrast, PnP-PDS (Unstable), which uses a denoiser that is not necessarily firmly nonexpansive, shows inconsistent behavior.
In Fig. 5, we provide the visual results for deblurring at $\sigma=0.01$, accompanied by PSNR (left) and SSIM (right). The images obtained by PnP-FBS and TV are overly-smoothed, resulting in a loss of fine detail. For the PnP methods without theoretical convergence guarantees, namely PnP-PDS (Unstable) and DRUNet, strong artifacts were generated because of the unstable behavior of the algorithm (see for example the area above the glove and the texture on the ground in the image, respectively). While RED achieves high PSNR and SSIM, there are some areas where the brightness appears unnaturally altered compared to the original image. The proposed method achieves the highest PSNR and SSIM for this image, presenting the most natural appearance.
Remark 3 (Comparison with Clamping of Denoiser Output).
An alternative approach to enforcing the box constraint is to clamp the denoiser output to [0, 1]01[0,\,1][ 0 , 1 ] after each iteration. This simple post-processing step can help stabilize the algorithm, and we evaluated its effect in the context of PnP-PDS without an explicit box constraint. Empirically, the performance was nearly identical to that of the proposed method across all settings in this section. However, the proposed formulation explicitly incorporates the box constraint into the monotone inclusion formulation. This allows for a clearer characterization of the optimality condition and improves the interpretability of the solution.
IV-C Deblurring/Inpainting under Poisson Noise
Scaling coefficient $\eta$ $50$ $100$ $200$
Deblurring $0.00125$ $0.00125$ $0.00100$
Inpainting $0.00075$ $0.00050$ $0.00050$
TABLE V: The Values of $\lambda$ Used in Each Setting for the Proposed Method.
Deblurring Inpainting
Scaling coefficient $\eta$ $50$ $100$ $200$ $50$ $100$ $200$
PnP-ADMM [24] $25.16$ $26.28$ $26.55$ $11.45$ $27.87$ $28.10$
PnP-PDS (Unstable) [25] $23.98$ $25.18$ $26.51$ $\mathbf{29.75}$ $\mathbf{31.90}$ $\mathbf{33.43}$
PnP-PDS (w/o a box const.) [26] $\mathbf{25.17}$ $\underline{26.42}$ $\mathbf{27.77}$ $\mathbf{29.75}$ $\underline{31.21}$ $\underline{32.74}$
BPG [50] $21.76$ $26.36$ $27.45$ $11.84$ $12.06$ $12.18$
RED [44] $25.13$ $26.39$ $27.72$ $29.71$ $31.19$ $32.69$
PnP-PDS (Proposed) $\mathbf{25.17}$ $\mathbf{26.44}$ $\mathbf{27.77}$ $\mathbf{29.75}$ $\underline{31.21}$ $\underline{32.74}$
TABLE VI: The Average PSNR [dB] Values for Deblurring and Inpainting on the Three Images from Set3[17]. The Best Value is Highlighted in Bold, and the Second Best Value is Underlined.
Deblurring Inpainting
Scaling coefficient $\eta$ $50$ $100$ $200$ $50$ $100$ $200$
PnP-ADMM [24] $1.692$ $2.224$ $2.898$ $1.203$ $0.6317$ $0.9953$
PnP-PDS (Proposed) $0.6523$ $0.6465$ $0.6529$ $0.4005$ $0.4030$ $0.4039$
TABLE VII: The Average CPU Time [s] per Iteration for Deblurring and Inpainting on the Three Images from Set3[17].
Figure 6: The evolution of PSNR [dB] and $c_{n}$ of PnP-PDS (Unstable) and the proposed method over iterations for deblurring at $\eta=100$. Each plot in the graph corresponds to each image extracted from Set3[17].
(a) (PSNR, SSIM)
Ground truth
(b) (19.26, 0.2653)19.260.2653(19.26,\,0.2653)( 19.26 , 0.2653 )
Observation
(c) $(25.17,\,\underline{0.7743})$
PnP-ADMM
(d) (24.05, 0.7618)24.050.7618(24.05,\,0.7618)( 24.05 , 0.7618 )
PnP-PDS
(e) (25.31¯, $(\underline{25.31},\,0.7731)$
PnP-PDS
(w/o a box const.)
(f) (25.31¯, $(\underline{25.31},\,0.7678)$
RED
(g) (24,74, $(24,74,\,\mathbf{0.7829})$
BPG
(h) (25.33, $(\mathbf{25.33},\,0.7733)$
Proposed
Figure 7: Visual results of image restoration on Set3 for deblurring at $\eta=100$. The best results are indicated in bold, and the second-best results are underlined.
Let us now present the experiments on deblurring/inpainting under Poisson noise. We considered three different scaling coefficients $\eta$ for the Poisson noise level, i.e. $50$, $100$, and $200$. Smaller values of $\eta$ correspond to stronger noise, while larger values indicate weaker noise. We compare the performance of the proposed method with other state-of-the-art methods applicable to this scenario, specifically PnP-ADMM [24], PnP-PDS (Unstable) [25], PnP-PDS (w/o a box const.) [26], RED [44], and BPG [50]. For methods whose parameter settings are suggested by the authors, we followed their recommendations. For the others, we performed linear searches over the balancing parameters between the data-fidelity and regularization terms. In particular, the best $\lambda$ in (35) values in the proposed method are summarized in Table V.
The average PSNR values over three images are presented in Table VI. The proposed method shows high restoration performance across all settings and outperforms the other state-of-the-art methods with higher PSNR values.
We investigate the computational efficiency of the proposed method, particularly in comparison with PnP-ADMM. Table VII provides the CPU times per iteration for PnP-ADMM and the proposed method. As discussed in Section III-C2, we need to perform inner iterations for PnP-ADMM, leading to relatively high computational costs. The proposed method reduces CPU times by half compared to PnP-ADMM in all settings, since it solves the problem in (35) without inner iterations.
To evaluate the stability, we present PSNR and the update rate $c_{n}$ at each iteration in the proposed method and PnP-PDS (Unstable) in Fig. 6. For PnP-PDS (Unstable), $c_{n}$ exhibits inconsistently depending on the image. In contrast, the proposed method shows a consistent decrease in $c_{n}$ over iterations, along with stable PSNR values, indicating the stable behavior of the proposed method.
Finally, we present the visual results of the deblurring at $\eta=100$ with PSNR (left) and SSIM (right) in Fig 7. PnP-PDS (Unstable) loses the details of the original images, while PnP-ADMM, RED, and PnP-PDS (w/o a box const.) show slight changes in brightness or fine shapes. Although BPG preserves the structure of the image well, it also produced several visually unpleasant artifacts (e.g., above the man’s shoulder and head). In contrast, the proposed method preserves fine details, resulting in an image that appears visually natural.
IV-D Influence of the Prior on the Proposed Method
Figure 8: The influence of the variation in $\sigma_{J}$ on the restoration performance for the proposed method. The plotted values are the average of seven images from ImageNet [65].
Finally, we study how the choice of prior influences the performance of the proposed method. As a representative case, we focus on how the noise level $\sigma_{J}$ assumed during the training of the Gaussian denoiser affects reconstruction performance of deblurring/inpainting under Gaussian noise.
[…]
The performance comparison is shown in Fig. 8. The noise level $\sigma$ of the observation was selected from seven values between $0.005$ and $0.1$. For each fixed $\sigma$, the proposed method exhibits nearly identical performance across all values of $\sigma_{J}$, with the maximum PSNR and SSIM differences being $1.215$ dB (in deblurring at $\sigma=0.005$) and $0.021$ (in deblurring at $\sigma=0.025$), respectively. These results indicate that the proposed method can effectively leverage the prior of denoisers trained with a wide range of $\sigma_{J}$. Furthermore, it achieves similarly competitive performance even when using a denoiser trained with randomly varying noise levels $\sigma_{J}\in[0,0.1]$, suggesting that its regularization effect is not highly sensitive to the specific training noise level in the proposed method.
Finally, we note that the consistent performance regardless of $\sigma_{J}$ and $\sigma$ can be attributed to expressing the data-fidelity as a constraint. By formulating it as a constraint, the consistency between the reconstruction and the observation can be explicitly controlled via the parameter $\varepsilon$. This helps prevent overly aggressive denoising that could arise when $\sigma_{J}$ is much larger than $\sigma$.
V Conclusion
In this paper, we have proposed a general PnP-PDS method with a theoretical convergence guarantee under realistic assumptions in real-world settings, supported by extensive experimental results. First, we established the convergent PnP-PDS using a firmly nonexpansive denoiser and theoretically characterized its solution set. Then, we showed that it efficiently solves various image restoration problems involving nonsmooth data-fidelity terms and additional hard constraints without the need to compute matrix inversions or perform inner iterations. Finally, through numerical experiments on deblurring/inpainting under Gaussian noise and deblurring/inpainting under Poisson noise, we demonstrated that the proposed PnP-PDS outperforms existing methods with stable convergence behavior. We believe that the proposed PnP-PDS holds great potential for a wide range of applications including biomedical imaging, astronomical imaging, and remote sensing.
[Proofs of Proposition 3.1 and Proposition 3.2] We start by introducing several key facts, followed by the proofs.
Fact 1 ([13, Lemma 4.4]).
Let $B_{1}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximally monotone operator, and $B_{2}:\mathcal{H}\to\mathcal{H}$ be a $\kappa$-cocoercive operator for some $\kappa\in(0,\,\infty)$. Suppose that $\mathrm{zer}(B_{1}+B_{2})$ is nonempty, and the following conditions hold:
  1. $(i)$
[…]
  2. $(ii)$
    $\rho_{n}\in\left(0,\,\delta\right)$,
  3. $(iii)$
    $\sum_{n\in\mathbb{N}}\rho_{n}(\delta-\rho_{n})=\infty$,
where $\delta:=2-\gamma\,(2\kappa)^{-1}$. Consider the sequence $\{\mathbf{s}_{n}\}_{n\in\mathbb{N}}$ generated by the following update:
$\begin{split}\mathbf{s}_{n+1}=\rho_{n}((\mathrm{Id}+\gamma B_{1})^{-1}(\mathrm{Id}-\gamma B_{2})(\mathbf{s}_{n})&)+\\ &(1-\rho_{n})\mathbf{s}_{n}.\end{split}$ (46)
The sequence $\{\mathbf{s}_{n}\}_{n\in\mathbb{N}}$ generated by (46) weakly converges to $\hat{\mathbf{s}}$ s.t. $\hat{\mathbf{s}}\in\mathrm{zer}(B_{1}+B_{2})$.
Fact 2 ([13, Lemma 4.2]).
Let $B:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximally monotone operator. Suppose that $\mathrm{zer}(B)$ is nonempty, and the following conditions hold:
  1. $(i)$
[…]
  2. $(ii)$
    $\sum_{n\in\mathbb{N}}\rho_{n}(2-\rho_{n})=\infty$.
Consider the sequence $\{\mathbf{s}_{n}\}_{n\in\mathbb{N}}$ generated by the following update:
$)\displaystyle\mathbf{s}_{n+1}=\rho_{n}((\mathrm{Id}+B)^{-1}(\mathbf{s}_{n})$$)+(1-\rho_{n})\mathbf{s}_{n}.$ (47)
The sequence $\{\mathbf{s}_{n}\}_{n\in\mathbb{N}}$ generated by (47) weakly converges to $\hat{\mathbf{s}}$ s.t. $\hat{\mathbf{s}}\in\mathrm{zer}(B)$.
Fact 3 ([13, Lemma 4.5]).
Let $J:\mathcal{H}\to\mathbb{R}$ be a convex differentiable function. If $\kappa\nabla J$ is nonexpansive for some $\kappa\in(0,\,\infty)$, then $\kappa\nabla J$ becomes firmly nonexpansive.
Proof of Proposition 3.1.
We follow the proof structure as in [13, Theorem 3.1 for Algorithm 3.1]. Consider the real Hilbert space $\mathcal{Z}_{I}=\mathcal{X}\times\mathcal{Y}$ equipped with the inner product defined as $\left\langle\mathbf{z},\,\mathbf{z^{\prime}}\right\rangle_{I}:=\left\langle\mathbf{x},\,\mathbf{x^{\prime}}\right\rangle+\left\langle\mathbf{y},\,\mathbf{y^{\prime}}\right\rangle$ where $\mathbf{z}=(\mathbf{x},\,\mathbf{y}),\,\mathbf{z^{\prime}}=(\mathbf{x^{\prime}},\,\mathbf{y^{\prime}})$. Now, we define $P$ as
$P:\begin{pmatrix}\mathbf{x}\\ \mathbf{y}\end{pmatrix}\to\begin{pmatrix}\mathrm{Id}/\gamma_{1}&-\mathbf{L}^{*}\\ -\mathbf{L}&\mathrm{Id}/\gamma_{2}\end{pmatrix}\begin{pmatrix}\mathbf{x}\\ \mathbf{y}\end{pmatrix}.$ (48)
Then, as shown in [13], $P$ is bounded, self-adjoint, and strictly positive. Hence, we obtain the real Hilbert space $\mathcal{Z}_{P}=\mathcal{X}\times\mathcal{Y}$ equipped with the following inner product:
$\left\langle\mathbf{z},\,\mathbf{z^{\prime}}\right\rangle_{P}:=\left\langle\mathbf{z},\,P\mathbf{z^{\prime}}\right\rangle_{I},$ (49)
Since convergence in $\mathcal{Z}_{I}$ and $\mathcal{Z}_{P}$ are equivalent [13, Theorem 3.1 for Algorithm 3.1], we show that the sequence generated by the iteration in (19) weakly converges to the solution to (20) in $\mathcal{Z}_{P}$.
For simple representation of Algorithm (19), let us define operators $T_{1}$ and $T_{2}$ as follows:
$T_{1}(\mathbf{z}_{n}):=\begin{pmatrix}\gamma_{1}^{-1}A_{J}(\mathbf{x}_{n})+\mathbf{L}^{*}\mathbf{y}_{n}\\ -\mathbf{L}\mathbf{x}_{n}+\partial h^{*}(\mathbf{y}_{n})\end{pmatrix},T_{2}(\mathbf{z}_{n}):=\begin{pmatrix}\nabla f(\mathbf{x}_{n})\\ 0\end{pmatrix},$ (50)
where $\mathbf{z}_{n}:=(\mathbf{x}_{n}^{\top},\,\mathbf{y}_{n}^{\top})^{\top}$. Then, we rewrite Algorithm (19) as
$\mathbf{z}_{n+1}=\rho_{n}(\mathrm{Id}+M_{1})^{-1}(\mathrm{Id}-M_{2})\mathbf{z}_{n}+(1-\rho_{n})\mathbf{z}_{n},$ (51)
where $M_{1}:=P^{-1}\circ T_{1}$ and $M_{2}:=P^{-1}\circ T_{2}$. Since the update in (51) has the same form as (46) with $\gamma=1$, it is sufficient to demonstrate that each assumption in Fact 1 is fulfilled. Hereunder, we define $\kappa$ as follows:
$\kappa:=\dfrac{1}{\beta}\left(\dfrac{1}{\gamma_{1}}-\gamma_{2}\|\mathbf{L}\|_{\mathrm{op}}^{2}\right).$ (52)
  • We see that the set-valued operator $A_{J}=J^{-1}-\mathrm{Id}$ is maximally monotone from [54, Proposition 23.8 (iii)]. Hence, $\gamma_{1}^{-1}A_{J}$ is also maximally monotone [54, Proposition 20.22]. The operator $\partial h^{*}$ is maximally monotone from [54, Theorem 20.48, Proposition 20.22, and Corollary 16.30]. Therefore, the operator $(\mathbf{x},\mathbf{y})\mapsto\gamma_{1}^{-1}A_{J}(\mathbf{x})\times\partial h^{*}(\mathbf{y})$ is maximally monotone in $\mathcal{Z}_{I}$ from [54, Proposition 20.23]. Furthermore, according to [54, Example 20.35], $(\mathbf{x},\mathbf{y})\mapsto\mathbf{L}^{*}\mathbf{y}\times-\mathbf{L}\mathbf{x}$ is maximally monotone and has a full domain. Therefore, $T_{1}$ becomes an maximally monotone operator on $\mathcal{Z}_{I}$ [54, Corollary 25.5 (i)]. Since $P$ is injective, $M_{1}=P^{-1}\circ T_{1}$ is maximally monotone on $\mathcal{Z}_{P}$.
  • From Fact 3 and the computation presented in [13, Theorem 3.1 for Algorithm 3.1], $M_{2}$ is cocoercive, i.e., $\kappa M_{2}$ is firmly nonexpansive in $\mathcal{Z}_{P}$.
  • We set $\gamma=1$. Since $10<2\kappa<$ 1 from $(i)$ of Proposition 3.1, the assumption $(i)$ of Fact 1 is fulfilled.
  • The assumption $(ii)$ of Fact 1 is fulfilled from (52).
  • The set of solutions to (20) coincides $\mathrm{zer}(T_{1}+T_{2})=\mathrm{zer}(M_{1}+M_{2})$, and this set is nonempty by the assumption.
Therefore, by applying Fact 1, we have that $\mathbf{z}_{n}$ weakly converges to $\mathbf{\hat{z}}\in\mathrm{zer}(T_{1}+T_{2})$ on $\mathcal{Z}_{P}$, which means $\{\mathbf{x}_{n},\mathbf{y}_{n}\}_{n\in\mathbb{N}}$ weakly converges to a solution to (20). ∎
Proof of Proposition 3.2.
Following the definition of $P$ as in Proposition 3.1, $P$ is bounded, linear, and self-adjoint. Since condition (i) does not include equality, $P$ is strictly positive. Therefore, same as the case in Proposition 3.1, we can define a Hilbert space $\mathcal{Z}_{P}$ equipped with the inner product $\langle\cdot|\cdot\rangle_{P}$. Since $f(\mathbf{x})=0$ , we have $T_{2}=0$, and the update in (51) becomes
$\mathbf{z}_{n+1}=\rho_{n}(\mathrm{Id}+M_{1})^{-1}\mathbf{z}_{n}+(1-\rho_{n})\mathbf{z}_{n}.$ (53)
The update in (53) is the same form as Fact 2 by setting $M=M_{1}$, so we need to show that the assumptions of Fact 2 are satisfied. Same as the proof of Proposition 3.1, $M_{1}$ becomes a maximally monotone operator, and from the assumption, $\mathrm{zer}(T_{1}+T_{2})=\mathrm{zer}(T_{1})=\mathrm{zer}(M_{1})$ is nonempty. Furthermore, the assumption on $\rho_{n}$ in Proposition 3.1 is the same as in Fact 2. Therefore, by Fact 2, $\mathbf{z_{n}}$ weakly converges to $\mathrm{zer}(T_{1}+T_{2})$, which means $\{\mathbf{x}_{n},\,\mathbf{y}_{n}\}$ weakly converges to the solution to (20). ∎